OpenStudy (anonymous):

A company's annual profit increases by $2,500 per year. In 2000, the profit of the company was $50,000. What was the total profit for 2007 if you assume the profit trend remained the same? A. $52,500 B. $65,000 C. $67,500 D. $70,000

OpenStudy (♪chibiterasu):

Basically we need to use: t = 50,000 + 2,500(7) First, multiply 7 by 2,500 since 7 is the different of years: 2,500 • 7 = 17,500 Now we have: t = 50,000 + 17,500 Now add: t = 67,500

OpenStudy (♪chibiterasu):

Did you get that?

OpenStudy (anonymous):

yea

OpenStudy (♪chibiterasu):

Lol your medal got me to 75 <3